Joe invests $2500 in an account that pays 4.75% per year. How much will he have if it is compounded monthly for 10 years?


1
Expert's answer
2021-03-10T15:02:12-0500

A0 = 2500 $, percentage per year is 4.75, but if it is compounded monthly, then the time 10 years is the same as 120 months. So the total amount of money will be A120 = A0 × (1 + 0.0475/12)120 = 2500 × (1 + 0.0475/12)120 = 4016.27$